raju is 19 yr younger than his cousin,after 5 yr their ages are in the ratio 2:3.what is their present age?
Answer by stanbon(75887)

raju is 19 yr younger than his cousin. After 5 yr their ages are in the ratio 2:3? What are their present ages?
-----------------
Equations:
r = c-19
----
(r+5)/(c+5) = 2/3
-----
cross-multiply to get:
3r+15 = 2c+10
---
Substitute for "r" and solve for "c":
3(c-19)+ 15 = 2c+10
3c-57+15 = 2c+10
c = 10+42
---
c = 52 (Cousin's age now)
---
Solve for "r":
r = c-19
r = 52-19
r = 33 (rugu's age now)
=========================
Cheers,
Stan H.
===========